40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1684     Entry time: Thu Jun 18 23:08:46 2009
Author: rob 
Type: Update 
Category: Locking 
Subject: spectrum 

Here's a noise spectrum of the RSE interferometer, in anti-spring mode, with RF readout.  I'd say the calibration is "loose."

I used the Buonanno & Chen modification of the KLMTV IFO transfer functions to model the DARM opto-mechanical response.  I just guessed at the quadrature, and normalized the optical gain at the frequency of the calibration line used (927Hz, not visible on the plot).

Attachment 1: DARMnoise_929352240.png  24 kB  | Hide | Hide all
DARMnoise_929352240.png
Attachment 2: DARMnoise_929352240.pdf  47 kB  | Hide | Hide all
DARMnoise_929352240.pdf
ELOG V3.1.3-